# High-dimensional dynamics in low-dimensional networks

### v1

@@ Abstract
Many networks that arise in nature and applications are effectively low-dimensional in the sense that their connectivity structure is dominated by a few dimensions. It is natural to expect that dynamics on such networks might also be low-dimensional. Indeed, recent results show that low-rank networks produce low-dimensional dynamics whenever the network is isolated from external perturbations or noise. However, networks in nature are rarely isolated. We show that recurrent networks with low-rank structure often produce high-dimensional dynamics in the presence of high-dimensional perturbations. Counter to intuition, dynamics in these networks are suppressed in directions that are aligned with the network’s low-rank structure, a phenomenon we term “low-rank suppression.” Our results clarify important, but counterintuitive relationships between a network’s connectivity structure and the structure of the dynamics it generates.
@@ Introduction
Recent work shows that many networks arising in nature and in applications are effectively low-rank in the sense that their connectivity matrices have a small number of larger singular values, while the remaining singular values are much smaller [1]. What are the implications of low-rank network structure on the dimensionality of network dynamics?
Several studies have shown that networks with low-rank structure tend to produce low-dimensional dynamics [1, 2, 3, 4]. However, these studies rely on specific assumptions about the network model. Some studies assume that the low-rank part of the network is weak [2, 4], others assume that the network is self-contained in the sense that it operates in the absence of internal noise or external perturbations or inputs [5, 1], while other work assumes that external perturbations are perfectly aligned with the network’s low-rank structure [3]. In reality, networks in nature are rarely isolated from perturbations, and these perturbations are not necessarily aligned with the network’s low-rank structure.
We use mathematical analysis and computer simulations to show that, in the presence of high-dimensional inputs or perturbations, many networks with low-dimensional structure produce high-dimensional dynamics. Moreover, perhaps counterintuitively, these networks suppress variability in directions that are aligned with their low-rank structure relative to random directions, an effect we call “low-rank suppression.” We show that many common structural features such as biased weights, modularity, and spatial connectivity structure promote low-rank suppression and high-dimensional dynamics. Finally, we demonstrate our conclusions in a real world epidemiological network with low-rank structure.
Our conclusions have important implications for the interpretation of low-dimensional structure in recurrent networks. In neuroscience, our results offer a mathematical explanation of the widely observed balance between excitatory and inhibitory synaptic currents [6, 7, 8, 9, 10, 11, 12] and also explain why neural populations generate high-dimensional responses to high-dimensional stimuli [13]. Beyond neuroscience, our results imply that randomly structured perturbations are often most effective at driving activity in networks with low-rank structure.
@@ Results
where ${\bf{x}}(t)$ is an external perturbation or input, ${\bf{z}}(t)$ is the network response, and we take $\tau=1$. The $N\times N$ connectivity matrix, $W$, takes the form
@@
then $W$ is “effectively low-rank” [1] in the sense that it has one large singular value near $|c|$ and the remaining singular values are bounded by $2\rho$ (Figure 1b). Similarly, $W$ has one large eigenvalue near $c$ and the remaining eigenvalues lie approximately within a circle of radius $\rho$ in the complex plane [14, 15]. Importantly, stability of the network dynamics requires that $\rho<1$ and $c<1$ [16]. Since low-rank structure requires $|c|\gg\rho$, stability requires $c<0$ when $\rho\sim\mathcal{O}(1)$. Here, we take $\rho=0.5$ and $c=-10$. We consider more general network structures later.
One might expect a network with connectivity $W$ to produce low-dimensional dynamics due to its low-rank structure. Indeed, a feedforward network with the same connectivity matrix produces approximately one-dimensional dynamics because it amplifies inputs aligned to ${\bf{u}}$ relative to other directions (Supplementary Figure S.1). We next demonstrate that this seemingly natural conclusion does not carry over to recurrent network dynamics.
Figure 1: Response properties of a recurrent network with rank-one structure. a) Schematic of model: The connectivity matrix, $W$, quantifies connections between nodes, ${\bf{z}}$, which receive external perturbations or input, ${\bf{x}}$. b) The singular values of $W$ have one dominant term, indicating approximate rank-one structure. c) The network response (${\bf{z}}(t)$, top) and its norm ($\|{\bf{z}}(t)\|$, bottom) given an input aligned to the low-rank structure of the network (${\bf{x}}_{\textrm{aligned}}$) and a random input (${\bf{x}}_{\textrm{rand}}$). d) The distribution of variance across principal components of a Gaussian stochastic input (${\bf{x}}(t)$; green) and the response (${\bf{z}}(t)$; blue). e,f) The network response (e) and input (f) projected onto the plane determined by ${\bf{u}}$ and a random vector, ${\bf{u}}_{\textrm{rand}}$, demonstrates low-rank suppression along ${\bf{u}}$. g) Local network input (purple) cancels with external input (green) to produce suppressed network responses (blue) in the direction of ${\bf{u}}$.
We first simulated the network with two different input patterns. The “aligned” input pattern was exactly aligned with the low-rank component of $W$,
@@
while the “random” input pattern is aligned with a random direction,
@@
which is generated identically to, but independently from ${\bf{u}}$. For large $N$, ${\bf{u}}$ and ${\bf{u}}_{\textrm{rand}}$ are approximately orthogonal. Intuitively, we might expect the network to respond more strongly to the aligned input than to the random input, as in a feedforward network. Perhaps surprisingly, we observed exactly the opposite (Figure 1c): The response to ${\bf{x}}_{\textrm{random}}$ was more than 11 times larger than the response to ${\bf{x}}_{\textrm{aligned}}$.
To demonstrate some consequences of this effect, we next provided a high-dimensional input, ${\bf{x}}(t)$, that varied smoothly in time. Specifically, each ${\bf{x}}_{j}(t)$ was an i.i.d. realization of a smooth, unbiased, stationary Gaussian process. This can model internal noise or irregular external input to the network.
Conventional wisdom and intuition might lead us to expect the network response to be low-dimensional and dominated by variability in the direction of ${\bf{u}}$, owing to the low-rank structure of $W$. In contrast, the variance explained by the principal components of ${\bf{z}}(t)$ decayed similarly to those of ${\bf{x}}(t)$ (Figure 1d), indicating that ${\bf{z}}(t)$ was high-dimensional like ${\bf{x}}(t)$. The only exception was the principal component that explained the least variance in ${\bf{z}}(t)$, which was much weaker than the other principal components (last blue dot in Figure 1d). Perhaps surprisingly, this weakest principal component direction was closely aligned to ${\bf{u}}$ (the angle between ${\bf{u}}$ and the corresponding principal component direction was less than $8^{\circ}$). Consistent with this finding, the variance of ${\bf{z}}(t)$ in the direction of ${\bf{u}}$ was more than 132 times smaller than the variance of ${\bf{z}}(t)$ along a random direction (Figure 1e) even though the variability of ${\bf{x}}(t)$ was similar in each direction (Figure 1f).
We propose the term “low-rank suppression” to refer to this phenomenon in which inputs aligned to dominant the low-rank structure of a network are suppressed by the network’s dynamics. In the absence of other directions that are amplified by the network, low-rank suppression leads to high-dimensional responses to high-dimensional perturbations, as in Figure 1d. In this manuscript, we derive conditions under which low-rank suppression and high-dimensional dynamics arise in low-rank networks. We then show that these conditions are met by many natural low-rank network structures.
The results in Figure 1 contrast with previous modeling work demonstrating low-dimensional dynamics in low-rank recurrent networks [2, 5, 3, 4, 1]. In some of these studies, external input or perturbations to the network are assumed to be absent [1, 5] or perfectly aligned with the low-rank component of the network connectivity [3]. In other studies [2, 4], the network is assumed to have only a weak low-rank component ($|c|\sim\rho$). None of these assumptions are consistent with the model considered in Figure 1. See Supplementary Text S.1 for a more detailed review of this previous work and its relation to ours.
A simple explanation for the low-rank suppression observed in Figure 1 is given by considering the steady state solution to Eq. (1),
@@
Convergence to the steady state requires that ${\bf{x}}(t)={\bf{x}}$ is static, but as long as ${\bf{x}}(t)$ varies more slowly than $\tau=1$, solutions approximately track the quasi-steady state given by Eq. (2). Because of the matrix inverse in Eq. (2), the large singular value of $W$ near $|c|$ produces a small singular value of $[I-W]^{-1}$ near $1/(1-|c|)\approx 1/|c|$ with left and right singular vectors near ${\bf{u}}$. Hence, the recurrent network suppresses inputs in the direction of ${\bf{u}}$. A related explanation is that the Jacobian matrix $J=W-I$ of Eq. (1) has a large, negative eigenvalue near $c-1<0$ with associated eigenvector near ${\bf{u}}$ so the dynamics in Eq. (1) are highly compressive along ${\bf{u}}$. These explanations depend on specific properties of the model considered here, but we later describe more general conditions under which low-rank suppression and high-dimensional dynamics occur.
@@
An interesting consequence of low-rank suppression is that external perturbations cancel nearly perfectly with network interactions in the direction of ${\bf{u}}$. More precisely, note that $W{\bf{z}}(t)$ in Eq. (1) can be interpreted as a vector of local inputs to each node whereas ${\bf{x}}(t)$ is external input, and ${\bf{z}}(t)$ is the network response. In the qausi-steady state, ${\bf{z}}\approx W{\bf{z}}+{\bf{x}}$. Under low-rank suppression, local input cancels nearly perfectly with external input ($W{\bf{z}}\approx-{\bf{x}}$) in the direction of ${\bf{u}}$ so that the response, ${\bf{z}}$, is weak in this direction (Figure 1g). This cancellation is a key feature of low-rank suppression.
@@
Conditions for low-rank suppression and high-dimensional dynamics
Above, we considered a simple example of a network with rank-one structure. In Supplementary Text S.2, we derive conditions for low-rank suppression and high-dimensional dynamics in a general class of models that includes all dynamics of the form $\tau d{\bf{z}}/dt={\bf{F}}({\bf{z}},{\bf{x}})$. Here, we summarize and interpret those results.
Low-rank suppression (as in Figure 1c) is a more general phenomenon than high-dimensional dynamics (as in Figure 1d). In Supplementary Text S.2, we show that low-rank suppression arises under two assumptions:
  1. 1.
    Assumptions on singular values. First, we assume that the network is strongly low-rank in the sense that the connectivity matrix $W$ has a small number of large singular values and the rest are much smaller (Figure 2a). Specifically, $W=W_{0}+W_{1}$ where $W_{0}$ has rank $r\ll N$ with large singular values ($\sigma_{W_{0}}\gg 1$) while $W_{1}$ has rank $N$ with small or moderate singular values ($\sigma_{W_{1}}\leq\mathcal{O}(1)$). “Weakly low-rank” networks for which $\sigma_{k}\sim\mathcal{O}(1)$ are studied in other work [2, 17, 4, 18, 19, 20] and do not necessarily produce low-rank suppression (see Supplementary Text S.1).
  2. 2.
    Assumptions on dynamics. In addition, our analysis relies on a linearization of the dynamics around a stable equilibrium. This requires that ${\bf{z}}(t)$ settles to a stable equilibrium whenever ${\bf{x}}(t)=0$, and that ${\bf{x}}(t)$ is sufficiently weak or the dynamics of ${\bf{z}}(t)$ are sufficiently linear. Nonlinear dynamics away from a stable equilibrium can produce any dynamics in principle [21, 22, 23] (see Supplementary Figures S.7 and S.8 for examples of unstable network states).
Under these two assumptions, low-rank suppression emerges in the sense that there are inputs, ${\bf{x}}$, approximately parallel to the column space of $U$, that produce much weaker responses than inputs at random directions (as in Figure 1c). The emergence of high-dimensional dynamics (as in Figure 1d) requires additional assumptions:
  1. 3.
    Assumptions on input perturbations. We assume that ${\bf{x}}(t)$ is a smooth, high-dimensional stochastic input. In simulations, we take each ${\bf{x}}_{j}(t)$ to be an i.i.d. Gaussian process. Low-dimensional inputs, ${\bf{x}}(t)$, can produce low-dimensional dynamics in networks that would otherwise produce high-dimensional dynamics (Supplementary Figure S.9).
  2. 4.
    Conditions on the alignment of singular vectors. Under the assumptions above, the emergence of high-dimensional dynamics depends on the structure of the low-rank part, $W_{0}$, of the connectivity. Specifically, let

### v2

@@ Abstract
Many networks in nature and applications have an approximate low-rank structure in the sense that their connectivity structure is dominated by a few dimensions. It is natural to expect that dynamics on such networks would also be low-dimensional. Indeed, theoretical results show that low-rank networks produce low-dimensional dynamics whenever the network is isolated from external perturbations or input. However, networks in nature are rarely isolated. Here, we study the dimensionality of dynamics in recurrent networks with low-dimensional structure driven by high-dimensional inputs or perturbations. We find that dynamics in such networks can be high- or low-dimensional and we derive precise conditions on the network structure under which dynamics are high-dimensional. In many low-rank networks, dynamics are suppressed in directions aligned with the network’s low-rank structure, a phenomenon we term “low-rank suppression.” We show that several low-rank network structures arising in nature satisfy the conditions for generating high-dimensional dynamics and low-rank suppression. Our results clarify important, but counterintuitive relationships between a recurrent network’s connectivity structure and the structure of the dynamics it generates.
@@ Introduction
Recent work shows that many networks arising in nature and applications have low-dimensional structure [1]. This observation raises an important question: What is the relationship between the dimensionality of a network’s structure and the dimensionality of the dynamics on the network?
In Neuroscience, for example, the implications of low-rank recurrent connectivity on neural dynamics is a topic of intense research [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. Theoretical studies demonstrating low-dimensional dynamics in low-rank neuronal network models are consistent with some neural recordings showing low-dimensional neural activity [13, 14, 15]. However, many of these recordings are made in the context of low-dimensional stimuli or behavior. A growing number of more recent studies have shown that neural activity can be high-dimensional, particularly in response to high-dimensional stimuli or behavior [16, 15, 17, 18, 19, 20]. These observations raise the important point that the dimensionality of dynamics on a network likely depends on the structure of the network’s external input, in addition to the network’s internal structure.
Some theoretical and computational studies demonstrating low-dimensional dynamics in low-rank networks focus on networks that are self-contained in the sense that they operate in the absence of noise or external inputs or perturbations [1, 5, 12]. Other work considers external input, but assumes that the input is perfectly aligned with the network’s low-dimensional structure [3]. In reality, networks in nature are rarely isolated from internal or external perturbations or inputs, and these perturbations are not necessarily aligned with the network’s low-dimensional structure. Other work [2, 4] considers the case of general external input, but the recurrent network is assumed to be weakly low-rank in the sense that the largest singular values are not asymptotically larger than the bulk of the singular values.
Here, we consider networks with high-dimensional external input and strongly low-rank structure, in which the largest few singular values are much larger than the bulk. We use mathematical analysis to derive precise conditions under which linearized dynamics in these networks produce high-dimensional dynamics. Perhaps counterintuitively, we find that these networks suppress dynamics in directions aligned to their low-dimensional structure, an effect we call “low-rank suppression.” We show that many common structural features such as biased weights, modularity, and spatial connectivity promote strongly low-rank structure, high-dimensional dynamics, and low-rank suppression. We also draw connections between low-rank suppression and the mathematical theory of balanced networks [21], extending previous work in this direction [3]. Finally, we demonstrate our conclusions in dynamics on a real epidemiological network.
Our conclusions have important implications for the interpretation of low-dimensional network structure. In neuroscience, our results can explain why neural populations generate high-dimensional responses to high-dimensional stimuli and tasks [16]. Our results also generalize and extend the theory of excitatory-inhibitory balance [21, 3], and amplification arising from breaks in this balance [22, 23, 24]. Beyond neuroscience, our results imply that perturbations misaligned to low-rank network structure are most effective at driving responses. This counterintuitive observation could be used to design more effective interventions to epidemiological, biological, social, and other networks.
@@ Results
where ${\bf{x}}(t)$ is an external perturbation or input, ${\bf{z}}(t)$ is the network response, and $\tau=1$. The $N\times N$ connectivity matrix, $W$, takes the form
@@
then $W$ is “effectively low-rank” [1] in the sense that it has one large singular value near $|c|$ and the remaining singular values are much smaller (bounded by $2\rho$; Figure 1b). Similarly, $W$ has one large eigenvalue near $c$ and the remaining eigenvalues lie approximately within a circle of radius $\rho$ in the complex plane (Figure 1c) [25, 26].
Importantly, stability of the network dynamics requires that $\rho<1$ and $c<1$ [27]. If the dynamics were unstable, then $\|{\bf{z}}(t)\|$ would grow exponentially toward $\infty$, so we only consider parameter regimes with stable dynamics. Since strong low-rank structure also requires $|c|\gg\rho$, stability requires $c<0$ when $\rho=\mathcal{O}(1)$. In simulations here, we take $\rho=0.5$ and c=10c=-10. We consider more general network structures later.
We first provided a high-dimensional input, ${\bf{x}}(t)$. Specifically, for our first simulation, each ${\bf{x}}_{j}(t)$ was an i.i.d., smooth, stationary Gaussian process, which models internal noise, external perturbations, or external input. Conventional wisdom might lead us to expect a low-dimensional network response dominated by variability in the direction of $\bf{u}$}. Indeed, a feedforward network with the same connectivity matrix and same input produces approximately one-dimensional dynamics because it amplifies inputs aligned to $\bf{u}$} (Supplementary Figure S.1). However, this conclusion does not necessarily carry over to recurrent networks.
In the recurrent network, the variance explained by the principal components of ${\bf{z}}(t)$ decayed similarly to those of ${\bf{x}}(t)$ (Figure 1d), indicating that ${\bf{z}}(t)$ was high-dimensional like ${\bf{x}}(t)$. The only exception was the principal component that explained the least variance in ${\bf{z}}(t)$, which was much weaker than the other principal components (last blue dot in Figure 1d). Perhaps surprisingly, this weakest principal component direction was closely aligned to $\bf{u}$} (the angle was less than $8^{\circ}$). Consistent with this finding, the variance of ${\bf{z}}(t)$ in the direction of $\bf{u}$} was more than 132 times smaller than the variance of ${\bf{z}}(t)$ along a random direction (Figure 1e) even though the variability of ${\bf{x}}(t)$ was similar in each direction (Figure 1f).
Figure 1: Response properties of a recurrent network with rank-one structure. a) Schematic of model: The connectivity matrix, $W$, quantifies connections between nodes, $\bf{z}$}, which receive external perturbations or input, $\bf{x}$}. b) The singular values of $W$ have one dominant term, indicating approximate rank-one structure. c) The eigenvalues of $W$ have a dominant, negative term. d) The distribution of variance across principal components of a Gaussian stochastic input (${\bf{x}}(t)$; green) and the response (${\bf{z}}(t)$; blue). e,f) The network response (e) and input (f) projected onto the plane determined by $\bf{u}$} and a random vector, ${\bf{u}}_{\textrm{rand}}$, demonstrates low-rank suppression along $\bf{u}$}. g) The network response (${\bf{z}}(t)$, top) and its norm ($\|{\bf{z}}(t)\|$, bottom) given an input aligned to the low-rank structure of the network (${\bf{x}}_{\textrm{aligned}}$) and a random input (${\bf{x}}_{\textrm{rand}}$). h) Local network input (purple) cancels with external input (green) to produce suppressed network responses (blue) in the direction of $\bf{u}$}.
To better understand these results, we next simulated the network with two different external input patterns. One pattern was aligned with the low-dimensional structure of $W$,
@@
while the other had a random direction,
@@
which was generated identically to, but independently from $\bf{u}$}. Intuitively, we might expect the network to respond more strongly to the aligned input than to the random input, as in a feedforward network (Supplementary Figure S.1). In reality, we observed exactly the opposite (Figure 1g): The response to ${\bf{x}}_{\textrm{random}}$ was more than 11 times larger than the response to ${\bf{x}}_{\textrm{aligned}}$.
We use the term “low-rank suppression” to refer to this phenomenon in which inputs aligned to the low-rank structure of a network are suppressed by the network’s dynamics. In the absence of other directions that are amplified by the network, low-rank suppression leads to high-dimensional responses to high-dimensional inputs (as in Figure 1d).
The results in Figure 1 contrast with some previous modeling work demonstrating low-dimensional dynamics in low-rank recurrent networks [2, 5, 3, 4, 1]. In Supplementary Text S.1, we provide a detailed review of this previous work and its relation to ours.
A simple explanation for the low-rank suppression observed in Figure 1 is given by considering the steady-state-solution to Eq. (1),
@@
Convergence to the steady-state requires that ${\bf{x}}(t)={\bf{x}}$ is static, but as long as ${\bf{x}}(t)$ varies more slowly than $\tau=1$, solutions approximately track the quasi-steady state given by Eq. (2). Because of the matrix inverse in Eq. (2), the large singular value of $W$ near $|c|$ produces a small singular value of $I-W]^{-1}$ near $1/(1-|c|)\approx 1/|c|$ with left and right singular vectors near $\bf{u}$}. Hence, the recurrent network suppresses inputs in the direction of $\bf{u}$}. A related explanation is that the Jacobian matrix J=WIJ=W-I of Eq. (1) has a large, negative eigenvalue near c1<0c-1<0 with associated eigenvector near $\bf{u}$} so the dynamics in Eq. (1) are highly compressive along $\bf{u}$}. This second explanation relies on the assumption that $W_{0}$ is a symmetric matrix (or at least a normal matrix) but we will next consider a case in which $W_{0}$ is non-normal and therefore non-symmetric.
@@
An interesting consequence of low-rank suppression is that external perturbations cancel nearly perfectly with recurrent inputs in the direction of $\bf{u}$}. More precisely, note that $W{\bf{z}}(t)$ in Eq. (1) can be interpreted as a vector of internal input to each node, whereas ${\bf{x}}(t)$ is external input and ${\bf{z}}(t)$ is the network response. In the qausi-steady state,
${\bf{z}}\approx W{\bf{z}}+{\bf{x}}.$ (3)
In other words, $\bf{z}$} tracks the sum of internal and external inputs in the quasi-steady state.
Since $W$ has a large singular value with left- and right singular values aligned to $\bf{u}$}, multiplication by $W$ amplifies the direction $\bf{u}$}. Therefore, the internal input, $W{\bf{z}}$, is large in the direction of $\bf{u}$} whenever $\bf{z}$} is moderate in the direction of $\bf{u}$}. In other words, $|{\bf{u}}\cdot W{\bf{z}}|\gg|{\bf{u}}\cdot{\bf{z}}|$. This fact might appear to present a paradox because the direction $\bf{u}$} is amplified in the product $W{\bf{z}}$, but $W{\bf{z}}$ is one component of $\bf{z}$} (Eq. (3)) and the direction $\bf{u}$} is suppressed in $\bf{z}$}.
This apparent paradox is resolved by a cancellation between ${\bf{W}}{\bf{z}}$ and $\bf{x}$} in the direction of $\bf{z}$}. Specifically, under low-rank suppression, internal input cancels nearly perfectly with external input in the direction of $\bf{u}$} (${\bf{u}}\cdot W{\bf{z}}\approx-{\bf{u}}\cdot{\bf{x}}$) so that the response, $\bf{z}$}, is weak in the direction of $\bf{u}$} (Figure 1h). We refer to this effect as “low-rank cancellation,” which is closely related to the theory of excitatory-inhibitory balance in neural circuits [21, 28] as we will show later.
@@
Low-dimensional dynamics and low-rank suppression in a non-normal network with rank-one structure
Our observation of low-rank suppression in the example above might seem unsurprising due to the presence of a large, negative eigenvalue. Likewise, our observation of high-dimensional dynamics might seem like a simple consequence of high-dimensional input. We next address these counterpoints with an example in which $W_{0}$ is highly non-normal. In this example, low-rank suppression arises in the absence of a large negative eigenvalue and the network’s response is low-dimensional even though the input is high-dimensional.
Figure 2: Response properties of a non-normal recurrent network with rank-one structure. Same as Figure 1 except the low-rank part of $W$ is non-normal. The network still exhibits low-rank suppression and cancellation (f, g, h) even in the absence of strong negative eigenvalues (c). The network exhibits low-dimensional dynamics (d), in contrast to Figure 1.
We specifically consider an example in which the model and all parameters are the same as in Figure 1 except that
$W_{0}=c{\bf{u}}{\bf{v}}^{T}$
where $\bf{u}$} and $\bf{v}$} are random, orthogonal vectors with $\|{\bf{u}}\|=\|{\bf{v}}\|=1$. This matrix is highly non-normal in the sense that its left- and right-singular vectors are orthogonal, and its non-zero eigenvalue does not have a simple relationship to its singular values.
The distribution of singular values of $W=W_{0}+W_{1}$ are identical to those from our previous example (Figure 2b; compare to Figure 1b), but the eigenvalues are different (Figure 2c; compare to Figure 1c). Specifically, the eigenvalues no longer have a dominant negative outlier. They are all $\mathcal{O}(1)$ in magnitude and clustered in a single bulk.
Repeating the simulations from above for this network, we found that the response to high-dimensional input was low-dimensional in the sense that variability was much greater in one direction than all others: Over 40% of the variance in ${\bf{z}}(t)$ was captured by a single principle component (first blue dot in Figure 2d), compared to less than 2% for any orthogonal direction (the remaining blue dots). This dominant direction was partially aligned to $\bf{u}$} (the angle between the dominant direction and $\bf{u}$} was $31^{\circ}$). Consistent with this observation, variability was $100$ times larger in the direction of $\bf{u}$} than in a random direction (Figure 2e). Hence, low-dimensional dynamics emerged even in the presence of high-dimensional input.
In addition, there was a direction in which variability was especially weak (last blue dot in Figure 2d), similar to our previous example. This last principal component direction, ${\bf{u}}_{N}$, was partially aligned to $\bf{v}$} (the angle between ${\bf{u}}_{N}$ and $\bf{v}$} was $13^{\circ}$). While the variability along $\bf{v}$} itself was not much smaller than a random directions (due to imperfect alignment between ${\bf{u}}_{N}$ and $\bf{v}$}), the variability along ${\bf{u}}_{N}$ was 190 times smaller than the variability in a random direction (Figure 2f). Also, low-rank cancellation was observed in the direction of ${\bf{u}}_{N}$ (Figure 2h).
Low-rank suppression was observed in the sense that an input in the direction of ${\bf{x}}_{\textrm{aligned}}={\bf{u}}+(W_{1}{\bf{v}}-{\bf{v}})/c$ produced a much weaker response than input in a random direction (Figure 2g). The choice of ${\bf{x}}_{\textrm{aligned}}$ will be discussed in a later section, but note that it is closely aligned to $\bf{u}$} whenever $c$ is large. In other words, inputs in the direction of ${\bf{x}}_{\textrm{aligned}}$, which is close to $\bf{u}$}, are suppressed by the network. Hence, low-rank suppression and low-rank cancellation arise even in the absence of a dominant negative eigenvalue.
To summarize our observations from the two examples above (Figures 1 and 2), both networks produced low-rank suppression and low-rank cancellation, but only the network with a non-normal connectivity matrix (Figure 2) produced low-dimensional dynamics. We next derive precise conditions for low-rank suppression and low-dimensional dynamics.
General conditions for low-rank suppression and high-dimensional dynamics
Above, we considered a simple network with rank-one structure. Here, we consider more general classes of networks, specify a list of assumptions that we make about the networks, and derive precise conditions for low-rank suppression and low-dimensional dynamics under these assumptions.
Assumption 1: Dynamics are linear or near a stable equilibrium.
We consider a general class of dynamics of the form
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+{\bf{F}}({\bf{z}},\bm{\eta})$ (4)
for some smooth vector field, ${\bf{F}}:{\mathbb{R}}^{N}\times{\mathbb{R}}^{N}\to{\mathbb{R}}^{N}$, where $\tau>0$ sets the timescale of the dynamics. Here, ${\bf{z}}(t)$ is an $N\times 1$ vector quantifying the state of $N$ interacting nodes and $\bm{\eta}(t)$ is an $N\times 1$ vector of external perturbations, inputs, and/or noise.
In general, the dependence of ${\bf{F}}({\bf{z}},\bm{\eta})$ on $\bf{z}$} defines interactions between nodes, but it is not immediately clear how to define a fixed network of interactions, $W$, from the dynamics in Eq. (4). When dynamics are linear, $\bf{F}$} can be written as ${\bf{F}}({\bf{z}},\bm{\eta})=W{\bf{z}}+{\bf{G}}(\bm{\eta})$, and we can take $W$ to represent the effective connectivity. In nonlinear systems, the effective connectivity between nodes can change with the state of the system. To resolve this difficulty in defining network structure when dynamics are nonlinear, we assume that dynamics represent a small deviation away from a stable equilibrium so that we can linearize them.
Specifically, we assume that there exists a time-constant forcing, $\bm{\eta}(t)=\bm{\eta}_{0}$, that produces a stable steady-state solution, ${\bf{z}}_{0}$. In other words, there exist $\bm{\eta}_{0}$ and ${\bf{z}}_{0}$ such that
${\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})={\bf{z}}_{0}.$
$J=-I+\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})$
has eigenvalues with strictly negative real part. Here, $I$ is the identity matrix and $\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})$ is the Jacobian matrix of $\bf{F}$} with respect to $\bf{z}$} evaluated at the fixed point. We then consider a small perturbation around this fixed point driven by a perturbation to the forcing term,
$\bm{\eta}_{p}(t)=\bm{\eta}_{0}+\epsilon\bm{\eta}(t).$
The response, ${\bf{z}}_{p}(t)$, of the network to the perturbed forcing term, $\bm{\eta}_{p}(t)$, can be written to linear order in $\epsilon$ as
${\bf{z}}_{p}(t)={\bf{z}}_{0}+\epsilon{\bf{z}}(t)+\mathcal{O}(\epsilon^{2}).$ (5)
The perturbation, ${\bf{z}}(t)$, obeys the linearized equation
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+W{\bf{z}}+{\bf{x}}(t)$ (6)
where
$W=\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})$ (7)
is interpreted as the effective connectivity matrix and
${\bf{x}}(t)=\partial_{\bm{\eta}}{\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})\bm{\eta}(t)$
is interpreted as the effective input. Our assumption that the equilibrium is stable implies that all eigenvalues of $W$ have real part less than $1$.
In many models, the perturbation is purely additive, ${\bf{F}}({\bf{z}},\bm{\eta})={\bf{H}}({\bf{z}})+\bm{\eta}$, so that the effective input is equal to the raw input, ${\bf{x}}=\bm{\eta}$. In other models, the raw input, $\bm{\eta}$, to the nonlinear system in Eq. (4) can have a different structure and dimensionality to the effective input, $\bf{x}$}, to the linearized system in Eq. (6), so care must be taken in interpreting our results. For illustrative purposes, we consider a contrived example of such a model in Supplementary Figure S.2.
Some models are parameterized in terms of a connectivity matrix that is not equal to the Jacobian matrix. For example, the system might be written as $d{\bf{z}}/dt={\bf{H}}(W{\bf{z}})+{\bf{x}}$ where $W$ is meant to quantify a connectivity matrix, but $\partial_{\bf{z}}{\bf{H}}(W{\bf{z}}_{0})\neq W$. In this case, our interpretation of the effective connectivity matrix is not consistent with the original parameterization, but they will often share structural properties such as their approximate dimensionality. We consider one such example from epidemiology at the end of the Results.
The assumption of linearized dynamics around a stable equilibrium is somewhat restrictive. However, nonlinear and unstable dynamics can produce any dynamics [29, 30], so it is difficult to make any general statements about the dimensionality and structure of dynamics in networks far from a stable equilibrium (see Supplementary Figures S.4 and S.5 for two examples of nonlinear, unstable dynamics). Moreover, the interpretation of network connectivity in nonlinear networks far from equilibrium is not clear, since effective interactions can change with network state. Previous work [1] considered models of the form $d{\bf{z}}/dt={\bf{G}}({\bf{y}},{\bf{z}})$ where ${\bf{y}}=W{\bf{z}}$ (note that this model lacks external input). This parameterization gives an explicit representation of the network structure, $W$, in nonlinear systems, and it allowed the authors to make statements about the dimensionality of dynamics away from equilibrium using global bounds on the Jacobian matrix, $\partial_{{\bf{y}}}{\bf{G}}({\bf{y}},{\bf{z}})$. We discuss the relationships between their analysis and ours in more depth in Supplementary Text S.1.
If a model is fully linear, i.e., if a model is already written in the form of Eq. (6), then we do not need to assume that dynamics remain close to the equilibrium, so inputs and perturbations need not be weak when dynamics are linear.
In summary, we assume that the network dynamics are either fully linear or, if they are nonlinear, they represent a small perturbation away from a stable equilibrium. This assumption allows us to focus our analysis on the linearized dynamics in Eq. (6).
Assumption 2: The connectivity matrix is strongly low rank in an asymptotic sense.
We now formalize an asymptotic notion of a strongly low-rank network. Specifically, we consider the limit of large network size, $N\to\infty$, and assume that $W$ has $r\sim\mathcal{O}(1)$ asymptotically large singular values,
$\sigma_{k}\gg 1,\;\;k=1,\ldots,r$
where $\sigma_{k}$ denotes the $k$th largest singular value of $W$. We assume that the remaining singular values are small or moderate in magnitude,
$\sigma_{j}\leq\mathcal{O}(1),\;\;j=r+1,\ldots,N.$
Later, we consider some examples in which these scaling laws arise naturally.
Matrices that satisfy the assumptions above can be decomposed as [1]
$W=W_{0}+W_{1}$
where $W_{0}$ has rank $r$ and asymptotically large singular values, while the singular values of $W_{1}$ are $\mathcal{O}(1)$ at most and $W_{1}$ can be full rank. Since $W_{0}$ has rank $r$, we can write its singular value decomposition as